A nail is fired into a block of wood at a velocity of 200 m/s. If the deceleration is 2 million m/s2 how far does the nail travel before it comes to a stop?


1
Expert's answer
2021-01-04T14:31:52-0500

Solution:

Assuming that deceleration keeps constant during the braking process, we can use one of the kinematics equations, as follows:

"v^2_f-v^2_o=2*a*(\Delta x)"

where "v_f" is the final velocity (0 in our case), "v_o" is the initial velocity (200 "\tfrac{m}{s}" ), decceleration (-2000000 "\tfrac{m}{s^2}" ), and Δx is the distance traveled.


"\dfrac{}{}"

"\Delta x=\dfrac{-v^2_o}{2*a}=\dfrac{-(200)^2}{2(-2000000)}"


"\Delta x=0.01\;m=1cm"


Answer:

That's a lot of deceleration! The nail would travel 1 cm.